# Multi-view Symbolic Regression

### v2

@@ Abstract.
Symbolic regression (SR) searches for analytical expressions representing the relationship between a set of explanatory and response variables. Current SR methods assume a single dataset extracted from a single experiment. Nevertheless, frequently, the researcher is confronted with multiple sets of results obtained from experiments conducted with different set-ups. Traditional SR methods may fail to find the underlying expression since the parameters of each experiment can be different. In this work we present Multi-View Symbolic Regression (MvSR), which takes into account multiple datasets simultaneously, mimicking experimental environments, and outputs a general parametric solution. This approach fits the evaluated expression to each independent dataset and returns a parametric family of functions $f(x;\theta)$ simultaneously capable of accurately fitting all datasets. We demonstrate the effectiveness of MvSR  using data generated from known expressions, as well as real-world data from astronomy, chemistry and economy, for which an a priori analytical expression is not available. Results show that MvSR obtains the correct expression more frequently and is robust to hyperparameters change. In real-world data, it is able to grasp the group behaviour, recovering known expressions from the literature as well as promising alternatives, thus enabling the use SR to a large range of experimental scenarios.
genetic programming, symbolic regression, regression
@@ 1. Introduction
The core of our modern scientific knowledge is based on careful production and analysis of experimental data. In a traditional scenario, the researcher’s task is to give meaning to measurable results, and analyze them in the light of an hypothesis. Frequently, the goal of this exercise is to find a mathematical description which can, at the same time, describe the recorded outcomes according to the state of the art of the field, and predict results from future similar experiments. Scientists became experts in the highly non-linear thought process required to interpret and translate scientific knowledge into suitable mathematical expressions. Due to the increase in data complexity, this task has also been approached through a machine learning perspective with the goal of automating the thought process.
@@
In many real scientific applications, the researcher is faced with different sources of data describing the same model but acquired from different viewpoints. In practice, this means that we have more data available to fit the model but, though they share the same functional structure, they may differ in parameter values. For example, consider data related to the contagion dynamics of a new virus collected from different populations. Even though the dynamics of the virus should be the same among all of the populations, the different aspects of a given population may influence the observed behavior in such a manner that all populations behave following a specific $f(x;\theta)$, but with different values of $\theta$. We argue that even if data of a single viewpoint (hereafter single-view) is abundant, multiple views can help constraining the search space of hypothesis.
In this paper, we introduce the idea of Multi-View Symbolic Regression (MvSR) that allows the practitioner to use the information from multiple sources and control the desired number of parameters to guide the search process to a parametric model that contains the correct number of parameters so it is neither too flexible (i.e., an universal approximator) nor too rigid (i.e., fitting only a single data source). This stimulates the discovery of scientific models that facilitates further analysis and interpretation of the phenomena. Specifically in this paper, we propose an adaptation to the Operon (Burlacu et al., 2020) package to calculate the fitness of a candidate solution by fitting this solution and calculating the base fitness individually on each data source and then returning an aggregated fitness of all sources. The experiments bring evidence that such procedure can successfully retrieve models capable of fitting all of the provided datasets.
@@ 2. Multi-View Symbolic Regression
To illustrate this concept, imagine that we have data from multiple epidemic events from different countries occasioned by the same pandemic. If we aggregate these data into a single dataset and fit a regression model, we will obtain a model that returns the expected contamination for a given time $t$. This expected value is calculated over the sum of all events and can be seen as explaining the average contagion event worldwide. On the other hand, if we have a model $f(t;\theta)$ fitted independently for each event, the parameters values will reveal different aspects for each of them. For example, fitting this model to data from different countries will describe different reproduction rates and thus inform about the characteristics of local populations (i.e., widespread use of mask, density of the population, etc.).
@@
However with MvSR  the objective becomes:
@@
    Allow parameters to be repeated.
@@
    Perform a minimization of the parameters independently for each dataset.
@@
    Penalise solutions based on the number of parameters used.
@@ 2.1. Implementation details
In this work we use Operon (Burlacu et al., 2020), a high-performance C++ framework supporting single and multi-objective GP with non-linear optimization of the parameters using the Levenberg-Marquardt algorithm  (Kommenda et al., 2020). This framework was reported to perform well with respect to the runtime and overall quality of the results (La Cava et al., 2021). The framework also has a Python module counterpart, called PyOperon, which presents all of the flexibility of the C++ version.
The implementation was adapted to support the creation of independent evaluators, each of which evaluates the symbolic expression to one of the inputted datasets. Each evaluator fits the expression to one dataset, such that the expression can have different adjusted parameters for each view, evaluate the loss function (i.e., mean squared error, log-likelihood), and then aggregate the losses using an aggregation function. This final value is then used during selection and reproduction stages to calculate the probability of survival.
@@ 3.1. Data generation
For the artificial data, we setup a series of challenging benchmarks based on standards from the SR literature. For this purpose we chose three generating functions:
@@
$f_{1}(x)$ is a third order polynomial function which constitutes a simple one dimensional case for which results can easily be visualised and interpreted. $f_{2}(x)$ and $f_{3}(x)$ are based on the Friedman functions (Friedman, 1991) for which free parameters have been added by replacing some constant values.
@@
Each of these benchmarks will contain a total of $4$ datasets (i.e., views) with the sample sizes fixed for each view. $f_{1}$ datasets consists of 20 points equally spaced on the interval $[-2,2]$. In the partial view case we keep the same number of points but we sample each dataset into evenly spaced domains $[-2,-1],[-1,0],[0,1]$ and $[1,2]$. $f_{2}$ and $f_{3}$ datasets consist of 100 points uniformly distributed on the intervals detailed in the scikit-learn package ${}^{,}$ [^2] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_friedman2.html.
@@
For each generative function, we create multiple datasets using different noise rates, $\left\{0.000,0.033,0.066,0.100\right\}$, in order to verify the robustness of this approach w.r.t. to noise. The noisy target is sampled from the distribution $\mathcal{N}(y,\sigma_{y}\sqrt{\frac{\tau}{1-\tau}})$, where $\tau$ is the noise rate.
@@ 3.2. Operon Hyperparameters and Post-processing
As mentioned in Section 2.1, we used an adapted version of pyOperon  [^3] https://github.com/heal-research/pyoperon/releases supporting the use of multiple datasets and aggregation function. For the following experiments we used the hyperparameters depicted in Table 2.
@@
Each one of the $176$ combinations of benchmark functions ($f_{1}$ partial domains, and $f_{1},f_{2},f_{3}$ multiple views), noise level and maximum size represents a single instance of our set of benchmarks. For each instance we ran pyOperon main SR module for each one of the four views independently as well as with the MvSR adaptation. Each experiment was repeated with $100$ different random seeds. After each run, the string representation of the best symbolic model is processed with Sympy to replace the numerical values with parameter variables and the corresponding expressions are stored for post-processing.
These expressions are individually refitted to the noiseless version of each dataset minimizing the least squares with the python package iminuit (Dembinski and et al., 2020). The final score is calculated using the mean squared error (MSE). Therefore, if the correct expression (or any equivalent) is generated the score will be 00, even for functions generated on noisy datasets. When larger trees are allowed, some runs may generate NaN values as output. In such case the value is replaced by infinity.
@@
population size
$1000$
number of evaluations
$100000000$
pool size
$5$
error metric
$MSE$
prob. cx
$1.0$
prob. mut.
$0.25$
max depth.
$10$
optim. iterations
$100$
aggregation function
max
operators
add, sub, mul, div, square, exp, sqrt, sin ($f_{2}$ only)
@@ 4. Artificial Benchmark Results
Figures 1 show heatmap plots of the obtained results for each benchmark function for every combination of noise level and maximum size. In these heatmaps, the colorscheme displays the average MSE of each combination, the lighter the color the better the result. Any values higher than 5 (including infinity) is depicted as the darkest color in order to keep the contrast in the visualization of the results. In each plot we show the worst and best single-view results (using the average of the plotted values as a choice criteria) and the MvSR results.
Figure 1. Heatmap of the average MSE of the tested combinations of noise and maximum expression size. Each row of heatmaps show results for the $f_{1}$, $f_{1}$ partial domains, $f_{2}$ and $f_{3}$ benchmarks, respectively. Columns represent the worst single-view (left), best single-view (center), and MvSR results (right). The colorbar represents the average MSE for that configuration, ranging from 00 (white) to a clipped value of $5$ dark blue. The clipping improves the comparison of small values.
From Figures 0(a), 0(b) and 0(c) we can see that for the noisy datasets and after a certain value of maximum size, both the single-view and MvSR behaves similarly reaching the correct solution. For the noiseless dataset, only MvSR is capable of finding the correct solution. This is expected since the single-view approach is submitted to a simplified version of the original expression and, thus, choose the simplest solution.
In Figures 0(d), 0(e) and 0(f) we can see a degradation of the results when using single-view with the noisy data. As the single-view has access to only one part of the polynomial curve, it can often fit the training data with a lower degree polynomial. This does not happen with the MvSR as it recovers a correct solution in almost every combination of noise and maximum size. The same behavior is displayed in Figures 0(g) to 0(l) where, in every combination, MvSR returns an optimal model (or near-optimal if the optimal model is not allowed by the maximum tree size) while the single-view approaches struggle with the lack of information and higher noise levels. We observe that MvSR is resilient against noise and successfully prevents overfitting even when the maximum size is larger than the original expression.
Another important question is whether MvSR returns a model with the correct number of parameters. A model with less than the optimal number of parameters is less flexible to fit the data from different sources and thus, has a tendency to underfit. Similarly, a model with more parameters than needed, can be too flexible and capable of fitting the noise imbued in the data, thus creating an overfitted model.
@@ 5. Scientific application
In this section we apply MvSR to real experimental datasets coming from 3 scientific fields, namely chemistry, finance and astrophysics. Such problems represent significantly harder challenge for SR algorithms as the data generally have no absolute “correct” generative model, is irregularly sampled and display non-Gaussian noise. However it constitute perfect testing grounds for MvSR, since one cannot know a priori if examples provide display the full or partial behavior of the best parametric model. Therefore, in such cases, the most conservative approach is to always use MvSR in order to build general laws.
@@ 5.1. Chemistry dataset
The Beer’s law is used to describe the properties of chemical species when $A\leq 1$. It states that the absorptive capacity of a dissolved substance is directly proportional to its concentration in a solution and is expressed as $A=\epsilon lc$ where A is the absorbance, $\epsilon$ is the molar extinction coefficient, l is the optical path length and c the concentration. A handful of alternative models have been proposed. They attempt to build a general Beer’s law capable of fitting a larger range of absorption. For example Bozdoğan (2022) suggested to use a quadratic polynomial equation to compensate the positive or negative deviations from the linearity. Yeh et al. (2023) proposed to extend law by adding two exponent parameters on $l$ and $c$, thus adding flexibility to describe the deviations. Both approaches extend the range of validity of the law but does not provide a solution general enough to properly characterize the absorption at any A value.
We propose to use MvSR on a set of measurements to find such a general parametric solution. In order to proceed, several wavelength scans were carried out using the Hitachi double-beam spectrophotometer UH3500 from 800 nm to 200 nm. Thus, four molecules were analyzed at various concentrations in dichloromethane, a commercial coumarin, two bodipy (Song et al., 2012)(Tran et al., 2023) and a porphyrin (tpp, 1967) which was previously synthesized in order to measure their absorbance A as a function of the solution’s concentration. Ideally the parametric model should recover their extinction coefficient $\epsilon$.
@@
The data acquired display high non-linearity behaviors, in particular the absorption tend to reach a plateau around $A=3$. Figure 3 presents the best fits along with the parameter values associated and the $R^{2}$ score. The extended Beer’s law proposed by MvSR carries ideal properties to fit the observations. Indeed, in the case where the linearity is respected, the $\mu$ parameter can be set to 0 and the equation simplifies into the original Beer’s law. In such case $\epsilon$ carry exactly the same information as the standard method. In the general case, $1/\mu$ characterizes the plateau that the absorption will reach at high concentrations. MvSR shows that an exponential transition between the linear evolution and the plateau provides an accurate fit to the data. We observe that all $R^{2}$ scores are close to 1, demonstrating that the law is fitted to the observations. Additionally, Figure 3 displays the classical Beer’s law fitted to the data points for which $A\leq 1$. We also observe that the lower the extinction coefficient is, the further apart the two models are at low absorption. This suggests a deviation from Beer-Lambert’s law at low $\epsilon$, and would require further investigation.
@@
In summary Equation 5 offers the possibility of computing molar extinction coefficients without being strictly restricted to the linear regime. The functional form proposed by MvSR contains two parameters and thus requires only a few data points to be constrained. It could be an easiest and quickest alternative for the determination of intrinsic properties of chemical species.

### v3

@@ Abstract.
Symbolic regression (SR) searches for analytical expressions representing the relationship between a set of explanatory and response variables. Current SR methods assume a single dataset extracted from a single experiment. Nevertheless, frequently, the researcher is confronted with multiple sets of results obtained from experiments conducted with different set-ups. Traditional SR methods may fail to find the underlying expression since the parameters of each experiment can be different. In this work we present Multi-View Symbolic Regression (MvSR), which takes into account multiple datasets simultaneously, mimicking experimental environments, and outputs a general parametric solution. This approach fits the evaluated expression to each independent dataset and returns a parametric family of functions $f(x;\theta)$ simultaneously capable of accurately fitting all datasets. We demonstrate the effectiveness of MvSR  using data generated from known expressions, as well as real-world data from astronomy, chemistry and economy, for which an a priori analytical expression is not available. Results show that MvSR obtains the correct expression more frequently and is robust to hyperparameters change. In real-world data, it is able to grasp the group behavior, recovering known expressions from the literature as well as promising alternatives, thus enabling the use SR to a large range of experimental scenarios.
genetic programming, symbolic regression, interpretability
@@ 1. Introduction
The core of our modern scientific knowledge is based on careful production and analysis of experimental data. In a traditional scenario, the researcher’s task is to give meaning to measurable results, and analyze them in the light of a hypothesis. Frequently, the goal of this exercise is to find a mathematical description which can, at the same time, describe the recorded outcomes according to the state of the art of the field, and predict results from future similar experiments. Scientists became experts in the highly non-linear thought process required to interpret and translate scientific knowledge into suitable mathematical expressions. Due to the increase in data complexity, this task has also been approached through a machine learning perspective with to automate the thought process.
@@
In many real scientific applications, the researcher is faced with different sources of data describing the same model but acquired from different viewpoints. In practice, this means that we have more data available to fit the model but, though they share the same functional structure, they may differ in parameter values. For example, consider data related to the contagion dynamics of a new virus collected from different populations. Even though the dynamics of the virus should be the same among all of the populations, the different aspects of a given population may influence the observed behavior in such a manner that all populations behave following a specific $f(x;\theta)$, but with different values of $\theta$. We argue that even if data of a single viewpoint (hereafter single-view) is abundant, multiple views can help to constrain the search space of hypothesis.
In this paper, we introduce the idea of Multi-View Symbolic Regression (MvSR) that allows the practitioner to use the information from multiple sources and control the desired number of parameters to guide the search process to a parametric model that contains the correct number of parameters so it is neither too flexible (i.e., a universal approximator) nor too rigid (i.e., fitting only a single data source). This stimulates the discovery of scientific models that facilitates the analysis and interpretation of the phenomena. Specifically, in this paper, we propose an adaptation to the Operon (Burlacu et al., 2020) package to calculate the fitness of a candidate solution by fitting this solution and calculating the base fitness individually on each data source and then returning an aggregated fitness of all sources. The experiments bring evidence that such procedure can successfully retrieve models capable of fitting all the provided datasets.
@@ 2. Multi-View Symbolic Regression
To illustrate this concept, imagine that we have data from multiple epidemic events from different countries caused by the same pandemic. If we aggregate these data into a single dataset and fit a regression model, we will obtain a model that returns the expected contamination for a given time $t$. This expected value is calculated over the sum of all events and can be seen as explaining the average contagion event worldwide. On the other hand, if we have a model $f(t;\theta)$ fitted independently for each event, the parameters values will reveal different aspects for each of them. For example, fitting this model to data from different countries will describe different reproduction rates and thus inform about the characteristics of local populations (i.e., widespread use of mask, density of the population, etc.).
@@
However, with MvSR  the objective becomes:
@@
    Allow reuse of parameters in the symbolic expression (i.e., $\theta_{0}$ can appear multiple times).
@@
    Optimize the parameters independently for each dataset.
@@
    Penalize solutions based on the number of parameters used.
@@ 2.1. Implementation details
In this work we use Operon (Burlacu et al., 2020), a high-performance C++ framework supporting single and multi-objective GP with non-linear optimization of the parameters using the Levenberg-Marquardt algorithm  (Kommenda et al., 2020). This framework was reported to perform well with respect to the runtime and overall quality of the results (La Cava et al., 2021). The framework also has a Python module counterpart, called PyOperon, which presents all the flexibility of the C++ version.
The implementation was adapted to support the creation of independent evaluators, each of which evaluates the symbolic expression to one of the inputted datasets. Each evaluator fits the expression to one dataset, such that the expression can have different adjusted parameters for each view, evaluate the loss function, and then aggregate the losses using an aggregation function. This value is used during selection and reproduction stages to calculate the probability of survival.
@@ 3.1. Data generation
For the artificial data, we set up a series of challenging benchmarks based on standards from the SR literature. For this purpose, we chose three generating functions:
@@
$f_{1}(x)$ is a third order polynomial function which constitutes a simple one dimensional case for which results can easily be visualized and interpreted. $f_{2}(x)$ and $f_{3}(x)$ are based on the Friedman functions (Friedman, 1991) for which free parameters have been added by replacing some constant values.
@@
Each of these benchmarks will contain a total of $4$ datasets (i.e., views) with the sample sizes fixed for each view. $f_{1}$ datasets consists of 20 points equally spaced on the interval $[-2,2]$. In the partial view case we keep the same number of points, but we sample each dataset into evenly spaced domains $[-2,-1],[-1,0],[0,1]$ and $[1,2]$. $f_{2}$ and $f_{3}$ datasets consist of 100 points uniformly distributed on the intervals detailed in the scikit-learn package  [^1] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.
@@
For each generative function, we create multiple datasets using different noise rates, $\left\{0.000,0.033,0.066,0.100\right\}$, to verify the robustness of this approach w.r.t. to noise. The noisy target is sampled from the distribution $\mathcal{N}(y,\sigma_{y}\sqrt{\frac{\tau}{1-\tau}})$, where $\tau$ is the noise rate.
@@ 3.2. Operon Hyperparameters and Post-processing
As mentioned in Section 2.1, we used an adapted version of pyOperon  [^2] https://github.com/heal-research/pyoperon/releases supporting the use of multiple datasets and aggregation function. For the following experiments, we used the hyperparameters depicted in Table 2.
@@
Each one of the $176$ combinations of functions (multiple views and, for $f_{1}$, partial domains), noise level and maximum size represents a single instance of our set of benchmarks. For each instance we ran pyOperon main SR module for each one of the four views independently as well as with the MvSR adaptation. Each experiment was repeated with $100$ different random seeds. After a run, the string representation of the best symbolic model is processed with Sympy to replace the numerical values with parameter variables and the corresponding expressions are stored for post-processing.
These expressions are individually refitted to the noiseless version of each dataset, minimizing the least squares with the python package iminuit (Dembinski and et al., 2020). The final score is calculated using the mean squared error (MSE). Therefore, if the correct expression (or any equivalent) is generated the score will be 00, even for functions generated on noisy datasets. When larger trees are allowed, some runs may generate NaN values as output. In such case, the value is replaced by infinity.
@@
population size $1000$
generations $1000$
pool size $5$
error metric $MSE$
prob. cx / mut. $1.0$ / $0.25$
max depth. $10$
optim. iterations $100$
agg. function max
operators $+$, --, $\times$, $\div$, ^ 2, exp, $\sqrt{}$, sin ($f_{2}$ only)
@@ 4. Artificial Benchmark Results
Figures 1 show heatmap plots of the results for each benchmark function for every combination of noise level and maximum size. The colorscheme displays the median MSE of each combination, the lighter the color the better the result. Any values higher than 5 (including infinity) is depicted as the darkest color in order to keep the contrast in the visualization of the results. In each plot, we show the worst and best single-view results (using the median of the plotted values as a choice criterion) and the MvSR results.
Figure 1. Heatmap of the median MSE of the tested combinations of noise and maximum expression size. Each row of heatmaps show results for the $f_{1}$, $f_{1}$ partial domains, $f_{2}$ and $f_{3}$ benchmarks, respectively. Columns represent the worst single-view (left), best single-view (center), and MvSR results (right). The colorbar represents the median MSE for that configuration, ranging from 00 (white) to a clipped value of $5$ dark blue. The clipping improves the comparison of small values.
Figures 1(a) to 1(c) show that for the noisy datasets and after a certain value of maximum size, both the single-view and MvSR behaves similarly reaching the correct solution. For the noiseless dataset, only MvSR is capable of finding the correct solution. This is expected since the single-view approach is submitted to a simplified version of the original expression and chooses the simplest solution.
In Figures 1(d), 1(e) and 1(f) we can see a degradation of the results when using single-view with the noisy data. As the single-view has access to only one part of the polynomial curve, it can often fit the training data with a lower degree polynomial. This does not happen with the MvSR as it recovers a correct solution in almost every combination of noise and maximum size. The same behavior is displayed in Figures 1(g) to 1(l) where, in every combination, MvSR returns an optimal model (or near-optimal if the optimal model is not allowed by the maximum tree size) while the single-view approaches struggle with the lack of information and higher noise levels. We observe that MvSR is resilient against noise and successfully prevents overfitting even when the maximum size is larger than the original expression.
Another important question is whether MvSR returns a model with the correct number of parameters. A model with less than the optimal number of parameters is less flexible to fit the data from different sources and, thus, has a tendency to underfit. Similarly, a model with more parameters than needed, can be too flexible and capable of fitting the noise imbued in the data, thus creating an overfitted model.
@@ 5. Scientific application
In this section we apply MvSR to real experimental datasets coming from three scientific fields, namely chemistry, finance, and astrophysics. Such problems represent a significantly harder challenge for SR algorithms as the data generally have no absolute “correct” generative model, are irregularly sampled and display non-Gaussian noise. However, it constitutes perfect testing grounds for MvSR, since one cannot know a priori if examples provide display the full or partial behavior of the best parametric model. Therefore, in such cases, the most conservative approach is to always use MvSR in order to build general laws.
@@ 5.1. Chemistry dataset
The Beer’s law is used to describe the properties of chemical species when $A\leq 1$. It states that the absorptive capacity of a dissolved substance is directly proportional to its concentration in a solution and is expressed as $A=\epsilon lc$ where A is the absorbance, $\epsilon$ is the molar extinction coefficient, l is the optical path length and c the concentration. A handful of alternative models have been proposed. They attempt to build a general Beer’s law capable of fitting a larger range of absorption. For example, Bozdoğan (2022) suggested using a quadratic polynomial equation to compensate the positive or negative deviations from the linearity. Yeh et al. (2023) proposed to extend law by adding two exponent parameters on $l$ and $c$, thus adding flexibility to describe the deviations. Both approaches extend the range of validity of the law but does not provide a solution general enough to properly characterize the absorption at any A value.
We propose to use MvSR on a set of measurements to find such a general parametric solution. In order to proceed, several wavelength scans were carried out using the Hitachi double-beam spectrophotometer UH3500 from 800 nm to 200 nm. Thus, four molecules were analyzed at various concentrations in dichloromethane, a commercial coumarin, two bodipy (Song et al., 2012)(Tran et al., 2023) and a porphyrin (tpp, 1967) which was previously synthesized in order to measure their absorbance A as a function of the solution’s concentration. Ideally, the parametric model should recover their extinction coefficient $\epsilon$.
@@
The data acquired display high non-linearity behaviors, in particular the absorption tend to reach a plateau around $A=3$. Figure 3 presents the best fits along with the parameter values associated and the $R^{2}$ score. The extended Beer’s law proposed by MvSR carries ideal properties to fit the observations. Indeed, in the case where the linearity is respected, the $\mu$ parameter can be set to 0 and the equation simplifies into the original Beer’s law. In such case, $\epsilon$ carry exactly the same information as the standard method. In the general case, $1/\mu$ characterizes the plateau that the absorption will reach at high concentrations. MvSR shows that an exponential transition between the linear evolution and the plateau provides an accurate fit to the data. We observe that all $R^{2}$ scores are close to 1, demonstrating that the law is fitted to the observations. Additionally, Figure 3 displays the classical Beer’s law fitted to the data points for which $A\leq 1$. We also observe that the lower the extinction coefficient is, the further apart the two models are at low absorption. This suggests a deviation from Beer-Lambert’s law at low $\epsilon$, and would require further investigation.
@@
In summary, Equation 5 offers the possibility of computing molar extinction coefficients without being strictly restricted to the linear regime. The functional form proposed by MvSR contains two parameters and thus requires only a few data points to be constrained. It could be an easiest and quickest alternative for the determination of intrinsic properties of chemical species.